# spectrum and local weak convergence of sparse random uniform hypergraphs

### v1

@@ Abstract.
Hypergraphs, Weak limit, Limiting spectral distribution
@@ 1. Introduction
A significant development in this direction is the notion of local convergence (also known as Benjamini–Schramm convergence) for sparse random graphs [4, 14, 1]. Specifically, when $G(n,p)$ satisfies $np=\lambda$ for a fixed $\lambda>0$, the graph rooted at a uniformly chosen vertex converges almost surely in a distribution called Galton–Watson measure (see [33, 7]). This result has been extended to higher-dimensional objects such as the Linial–Meshulam random simplicial complex in [2]. In the present work, we investigate the local weak limits of $k$-uniform random hypergraphs.
@@
The Erdős–Rényi random graph $G(n,p)$ admits a natural generalization to the $k$-uniform random hypergraphs. See [35, 36], among others. For $k\leq n$, given the vertex set $n]={1,2,\ldots,n}$, the hyperedges of a $k$-uniform random hypergraph are subsets drawn from $\mathcal{V}_{k}=\left[\binom{n}{k}\right$], the collection of all $k$-element subsets of $[n]$, independently with probability $p$.
For graphs, the local structure around a vertex is typically captured by a rooted graph centered at the vertex. In the context of $k$-uniform hypergraphs, the notion of a root can be extended beyond vertices to include $r$-cardinality subsets of the vertex set, where $1\leq r\leq k-1$. Such connections have been explored in the literature under various frameworks, including $r$-intersection graphs [45, 32], $r$-walks in hypergraphs [41], and other related notions. As an $r$-cardinality subset of the vertex set is assumed as a root. We refer to it as $r$-set rooted $k$-uniform hypergraph.
In this work, we show that the local structure of such hypergraphs converges to a random infinite $r$-set rooted hypergraph, which can be regarded as a generalization of both the Galton–Watson measure with Poisson offspring and the limiting object is similar to a $d$-block Galton–Watson measure with Poisson offspring distribution mentioned in [2], see Section 2.4 for precise definition. The local structure of $r$-set rooted $k$-uniform hypergraph is captured by the $r$-set rooted local weak limit of $k$-uniform hypergraphs. See Section 2.2 for a detailed description of $r$-set rooted local weak limit and associated topology. In 3.3, we show that the $r$-set rooted local weak limit of sparse $k$-uniform random hypergraphs tends to the $(\binom{k}{r}-1)$-block Galton–Watson measure with Poisson offspring. For $k=2$, and $r=1$, the 3.3 leads to the well known Benjamini-Schramm convergence proved by Benjamini and Schramm[14], and by Aldous and Steele[4] see 3.2. For the $r=1$ case, when each vertex of the random hypergraph serves as a root, the local weak limit of sparse random $k$-uniform hypergraphs tends to the (k1)(k-1)-block Galton–Watson measure with Poisson offspring, see 3.1. For r=k1r=k-1, the (k1)(k-1)-rooted $k$-uniform hypergraphs resembles the Linial-Meshulam model [39, 42]. For this case, also the limiting measure is (k1)(k-1)-block Galton–Watson measure with Poisson offspring, See 3.4. This result has been established in [2].
@@
Since the behavior of graph operators such as the adjacency matrix is governed by the local structure of the graph, this local convergence has direct implications for spectral properties. For a finite graph $G$, the adjacency matrix $A_{G}$ of $G$ is given by, for any complex-valued function $f$ on its vertex set,
@@
This type of spectral convergence has also been extended to random simplicial complexes, as shown in [2, Section 3]. In the present work, we extend this framework to random hypergraphs by leveraging our results on local convergence in the space of $r$-set rooted $k$-uniform random hypergraphs. In 3.5, we establish the limiting spectral distribution of the $r$-set adjacency matrices of $k$-uniform random hypergraphs. When $r=1$, this matrix generalizes several classical notions of hypergraph adjacency matrices (see [9, 10]), and its limiting spectral distribution is analyzed in 3.7. For r=k1r=k-1, the limiting spectral behavior aligns with that of the Linial–Meshulam model, as discussed in [2], and is presented in 3.8.
The rest of the paper is organized as follows. In Section 2, we recall all the essential definitions and notations to state our main results. The main results and their Corollaries are stated in Section 3. The proofs of 3.1 and 3.3 are given in Section 4 and Section 5 respectively. The proofs of both results proceed in two stages. Convergence in expectation is shown in Section 4.1 and Section 5.1, while almost sure convergence is established in Section 4.2 and Section 5.2. In Section 6 we provide the proofs of 3.5 and its Corollaries.
@@ 2. Preliminaries
In this section, we introduce essential definitions and fundamental concepts related to random graphs, their generalizations to hypergraphs, rooted graphs along with their associated topology, weak convergence, and the Galton-Watson measure. These notions serve as the theoretical framework for analyzing the weak convergence of sparse random hypergraphs to their limiting structures. We begin by recalling the Erdős–Rényi random graph [29] model. This serves as a foundation for extending the concept to uniform hypergraphs.
@@ 2.1. Random graphs and hypergraphs
Let $V\subseteq\mathbb{N}$ be a non-empty set and $E\subset\{(x,y)\;:\;x,y\in V,x\neq y\}$. A weighted graph $G$ is an order triple $G=(V,E,w)$, where $w:E\to\mathbb{R}$ is a real values function on $E$. A multigraph $G$ is a weighted graph with positive integer valued weights, that is, $w\in\mathbb{N}$. If $w\equiv 1$ then the $G$ is called a simple graph.
A hypergraph $H$ generalizes the concept of a graph by allowing hyperedges, which may connect more than two vertices. Formally, a hypergraph $H$ is an ordered pair $H=(V(H),E(H))$, where $V(H)$ is a non-empty set of vertices, and $E(H)$ is a collection of non-empty subsets of $V(H)$, called hyperedges. Given an integer $k\leq|V(H)|$, a hypergraph $H$ is called $k$-uniform if every hyperedge $e\in E(H)$ has cardinality $|e|=k$. Recall that $|.|$ denotes numbers of elements in the set.
@@
In random graphs and hypergraphs, the edges are random objects. For example, if we associate any $2$-element subset of the vertex set to a Bernoulli random variable, where success means this two-element subset forms an edge and failure means this two-element subset is not an edge. Similarly, for a $k$-uniform hypergraph, we associate any $k$-cardinality subset of the vertex set to a Bernoulli random variable.
2.1.1. Erdős–Rényi Graph :
Let $n$ be a natural number and let $p\in[0,1]$. The Erdős–Rényi random graph, denoted $\mathcal{G}(n,p)$, is a random graph on $n$ vertices where each edge is inclued with probability $p$ and independently [25], [5, Chapter-10]. In other words, the Erdős–Rényi graph $\mathcal{G}(n,p)$ is a probability space $(\mathbb{G}(n),2^{\mathbb{G}(n)},{\bf P})$, where $\mathbb{G}(n)$ denotes the collection of all graphs on $n$ vertices. The probability of a particular graph $G\in\mathbb{G}(n)$ is given by:
@@ 2.1.2. $k$-uniform random hypergraphs
The $k$-uniform random hypergraphs is a natural generalization of the Erdős–Rényi Graphs. For any integer $k\geq 2$, the $k$-uniform random hypergraph, denoted by $H(n,k,p)$, is a random hypergraph on $n$ vertices, where each $k$-set of $n]=\{1,\ldots,n\}$ is included as a hyperedge with probability $p$ and independently. More precisely, given integers $n\geq 2$ and $1\leq k\leq n$, the $k$-uniform random hypergraph $H(n,k,p)$ is the probability space $(\mathbb{H}^{k}_{n},2^{\mathbb{H}^{k}_{n}},{\bf P})$, where $\mathbb{H}^{k}_{n}$ is the collection of all $k$-uniform hypergraphs with the vertex set $[n]$, and $2^{\mathbb{H}^{k}_{n}}$ is its power set. That is, for $H\in\mathbb{H}^{k}_{n}$ with edge set $E(H)$, the probability measure is given by
@@
2.1.3. Linial Meshulam complex
Another most studied generalizations of the Erdős-Rényi graph in higher dimensions is the Linial-Meshulam complex, which will appear later in our discussion. This model was introduced by Linial and Meshulam [39] for dimension $2$ and by Wallach [43] for higher dimensions.
Let $V$ be a finite set and $\mathcal{P}(V)$ denote the set of all subsets of $V$. A collection $X\subset\mathcal{P}(V)$ is said to be a simplicial complex with vertex set $V$ if $\tau\in X$ and $\sigma\subset\tau$, then $\sigma\in X$. An element of the simplicial complex is called a simplex. The dimension of a simplex $\sigma$ is defined as $|\sigma|-1$, where $|\sigma|$ denotes the number of elements in $\sigma$. An element of dimension $j$ is called a $j$-simplex. The dimension of a non-empty simplicial complex is the highest dimension among its simplices.
For a given positive integer $k\geq 2$, the Linial-Meshulam complex of dimension $k$, denoted by $Y_{k}(n,p)$, is a random simplicial complex on $n$ vertices with a complete (k1)(k-1)-dimensional skeleton and each $k$-simplices included independently with probability $p\in[0,1]$. In particular, $k=2$, the Linial Meshulam complex of dimension $2$ is the Erdős-Rényi graph.
Observe that the Linial-Meshulam complex of dimension $k$ is a random hypergraph where all simplexes of dimension less than (k1)(k-1) are included as hyperedges and each $k$-dimensional simplex is included as a hyperedge with probability $p$ independently. However, it is not a $k$-uniform random hypergraph.
@@ 2.1.4. $r$-set weighted line graph
Figure 2. $H$ is a $3$-uniform hypergraph and $G_{2}(H)$ is the $2$-set line graph of $H$. Three hyperedges of $H$ correspond to three triangles, and the rest of the 22-sets are isolated vertices in this $2$-set line graph.
It is worth noting that the line graph of $Y_{k}(n,p)$ is defined in [2] as the graph $(\mathcal{V}_{k},E_{n})$, where $\mathcal{V}_{k}=[\binom{n}{k}$] is the set of all (k1)(k-1)-dimensional simplexes and $(\sigma,\sigma^{\prime})\in E_{n}$ if $\sigma\cup\sigma^{\prime}\in Y_{k}(n,p)$ for $\sigma,\sigma^{\prime}\in\mathcal{V}_{k}$. Observe that, the line graph of $Y_{k-1}(n,p)$ is the same as the (k1)(k-1)-set weighted line graph of $H(n,k,p)$.
@@ 2.3. Local weak convergence of hypergraphs
(4) $U_{r}(H)=U(G_{r}(H)),$
where $U_{r}(H)$ is the probability measure arising from a uniformly chosen root of $H$. In that case $U_{r}(H)$ is a probability measure on $\mathcal{P}(\mathbb{G}_{w}^{*})$. Given a sequence of hypergraphs $\{H_{n}\}_{n\geq 1}$ if $\{U_{r}(H_{n})\}$ converges weakly to $\nu$, that is,
@@
then $\nu$ is called the $r$-set local weak limit of the sequence of hypergraphs $\{H_{n}\}_{n\geq 1}$. We show that, $1\leq r\leq k-1$, the $r$-set local weak limit of the sequence $\{H_{n}\}$ with $H_{n}\in H(n,k,p)$ is the $d$-block Galton-Watson measure with $d=\binom{k}{r}-1$ and Poisson offspring distribution almost surely. We define the $d$-block Galton-Watson measure in the next subsection.
@@ 2.4.1. Galton-Watson measure
    The number of children of the root distributed as ${\bf P}_{X}$.
@@
    Each non-root vertex is reached via an edge from its parent. If such a vertex has total degree $k$ (including the parent edge), then its number of children is k1k-1. The distribution of its number of children is the shifted size-biased distribution $\hat{{\bf P}}_{X}$ given by
@@
For any natural number $d$, a $d$-block Galton-Watson tree is a random graph generated by the Galton-Watson branching process in which, instead of generating a single vertex, each vertex, as offspring, produces a set of $d$ vertices called a $d$-block. Thus, if an individual in the population has a total of $n$ offspring $d$-block, then it has a total of $nd$-number of offspring.
@@
    Conditioned on the number of children, the subgraphs rooted at each child ( that is, each member of child $d$-block) are independent and identically distributed copies of the $d$-block Galton–Watson graph.
@@
Now we discuss below about the rooted graphs that belongs to the support of $\nu_{d,\lambda}$. The vertex set of such a rooted graph can be embedded into the set
@@
[…]
@@
[…]
@@
Here $Z_{n}$ represents the number of individuals in the $n$-th generation, and $X_{i,n}$ denotes the number of offspring $d$-block of the $i$-th individual in that generation. The offspring of an individual $i$ in generation $n$ are labeled by
@@
(b) 22-set line graph of $3$-uniform hypergraph
@@
Figure 3. A 22-set rooted hypergraph, and its $2$-set line graph.
We refer the reader to [2, Section 4.2] for a detailed description of $d$-block Galton-Watson tree and the corresponding measure. Since $\nu_{d,\lambda}\in\mathcal{P}(\mathbb{G}^{*})$, that is a probability measure on the collection of all rooted graphs, $\nu_{d,\lambda}$ can be represented as a random rooted graph, we denote it as $(GW_{d},o)$.
For $d=\binom{k}{r}-1$, the $(\binom{k}{r}-1)$-block Galton–Watson tree $(GW_{d},o)$ admits a natural representation as the $r$-set line graph of an $r$-set rooted $k$-uniform hypergraph $(T_{k},S)$, where $|S|=r$. A root $o$ corresponds to an $r$-element subset $S$ of the vertex set $V(H)$. More generally, each vertex $v$ in $GW_{d}$ corresponds to some $r$-cardinality subset $S\subset V(H)$. Every $(\binom{k}{r}-1)$-block of offspring of $v$ arises from a hyperedge $e$ of $H$ that contains $S$. Since $e$ consists of $k$ vertices, it contains exactly $\binom{k}{r}$ distinct $r$-subsets. Among these, one is $S$ itself, while the remaining $\binom{k}{r}-1$ subsets correspond to the vertices in the offspring $(\binom{k}{r}-1)$-block of $v$. Thus, $(GW_{d},o)$ is the $r$-set line graph of an $r$-set rooted $k$-uniform hypergraph $(T_{k},S)$. A $3$-uniform hypergraph $H$ and its $2$-set underlying graph is a $2$-block Galton-Watson tree illustrated in Figure 3.
2.4.3. Structure of $(T_{k},S_{r})$.

### v2

@@ Abstract.
Hypergraphs; Random hypergraphs; Weak limit; Limiting spectral distribution
@@ 1. Introduction
A significant development in this direction is the notion of local convergence (also known as Benjamini–Schramm convergence) for sparse random graphs [4, 14, 1]. Specifically, when $G(n,p)$ satisfies $np=\lambda$ for a fixed $\lambda>0$, the graph rooted at a uniformly chosen vertex converges almost surely to a distribution called Galton–Watson measure (see [33, 7]). This result has been extended to higher-dimensional objects such as the Linial–Meshulam random simplicial complex in [2]. In the present work, we investigate the local weak limits of $k$-uniform random hypergraphs.
@@
The Erdős–Rényi random graph $G(n,p)$ admits a natural generalization to the $k$-uniform random hypergraphs. See [35, 36, 48], among others. For $k\leq n$, given the vertex set $n]={1,2,\ldots,n}$, the hyperedges of a $k$-uniform random hypergraph are subsets drawn from $\mathcal{V}_{k}=\left[\binom{n}{k}\right$], the collection of all $k$-element subsets of $[n]$, independently with probability $p$.
For graphs, the local structure around a vertex is typically captured by a rooted graph centered at the vertex. In the context of $k$-uniform hypergraphs, the notion of a root can be extended beyond vertices to include $r$-cardinality subsets of the vertex set, where $1\leq r\leq k-1$. Such connections have been explored in the literature under various frameworks, including $r$-intersection graphs [45, 32], $r$-walks in hypergraphs [41], and other related notions. We take an $r$-subset of the vertex set as the root and refer to the resulting object as an $r$-set rooted $k$-uniform hypergraph.
In this work, we show that the local structure of such hypergraphs converges to a random infinite $r$-set rooted hypergraph, which can be regarded as a generalization of both the Galton–Watson measure with Poisson offspring and the limiting object is similar to a $d$-block Galton–Watson measure with Poisson offspring distribution mentioned in [2], see Section 2.4 for precise definition. In particular, for the case $r=1$ the multi-type Galton–Watson hypertree studied in [48, 52]. The local structure of $r$-set rooted $k$-uniform hypergraph is captured by the $r$-set rooted local weak limit of $k$-uniform hypergraphs. See Section 2.2 for a detailed description of $r$-set rooted local weak limit and associated topology. In Theorem 3.3, we show that the $r$-set rooted local weak limit of sparse $k$-uniform random hypergraphs tends to the $(\binom{k}{r}-1)$-block Galton–Watson measure with Poisson offspring. For $k=2$ and $r=1$, the Theorem 3.3 leads to the well-known Benjamini-Schramm convergence, see Corollary 3.2. For the $r=1$ case, when each vertex of the random hypergraph serves as a root, the local weak limit of sparse random $k$-uniform hypergraphs tends to the (k1)(k-1)-block Galton–Watson measure with Poisson offspring, see Theorem 3.1. For r=k1r=k-1, the (k1)(k-1)-rooted $k$-uniform hypergraphs resembles the Linial-Meshulam model [39, 42]. In this case as well, the limiting measure is the (k1)(k-1)-block Galton–Watson measure with Poisson offspring (see Corollary 3.4). This result has been established in [2].
@@
Since the behavior of graph operators such as the adjacency matrix is governed by the local structure of the graph, this local convergence has direct implications for spectral properties. For a finite graph $G$, the adjacency matrix $A_{G}$ of $G$ acts on any complex-valued function $f$ on its vertex set by
@@
This type of spectral convergence has also been extended to random simplicial complexes, as shown in [2, Section 3]. In the present work, we extend this framework to random hypergraphs by leveraging our results on local convergence in the space of $r$-set rooted $k$-uniform random hypergraphs. In Theorem 3.5, we establish the limiting spectral distribution of the $r$-set adjacency matrices of $k$-uniform random hypergraphs. When $r=1$, this matrix generalizes several classical notions of hypergraph adjacency matrices (see [9, 10]), and its limiting spectral distribution is analyzed in Corollary 3.7. For r=k1r=k-1, the limiting spectral behavior aligns with that of the Linial–Meshulam model, as discussed in [2], and is presented in Corollary 3.8.
The rest of the paper is organized as follows. In Section 2, we recall all the essential definitions and notations to state our main results. The main results and their corollaries are stated in Section 3. The proofs of Theorem 3.1 and Theorem 3.3 are given in Section 4 and Section 5 respectively. The proofs of both results proceed in two stages. Convergence in expectation is shown in Section 4.1 and Section 5.1, while almost sure convergence is established in Proposition 4.2 and Section 5.2. In Section 6, we provide the proofs of Theorem 3.5 and its corollaries.
@@ 2. Preliminaries
We begin by setting out the basic concepts that will be used throughout: random graphs and hypergraphs, rooted graphs and their associated topology, weak convergence, and the Galton–Watson measure. These notions form the theoretical framework for analyzing the weak convergence of sparse random hypergraphs to their limiting structures. We begin by recalling the Erdős–Rényi random graph [29]. This serves as a foundation for extending the concept to uniform hypergraphs.
@@ 2.1. Random graphs and hypergraphs
Let $V\subseteq\mathbb{N}$ be a non-empty set and $E\subset\{\{x,y\}\;:\;x,y\in V,x\neq y\}$. A weighted graph $G$ is an ordered triple $G=(V,E,w)$, where $w:E\to\mathbb{R}$ is a real-valued function on $E$. A multigraph $G$ is a weighted graph with positive integer valued weights, that is, $w(E)\subseteq\mathbb{N}$. If $w\equiv 1$ then the $G$ is called a simple graph.
A hypergraph $H$ generalizes the concept of a graph by allowing hyperedges, which may connect more than two vertices. Formally, a hypergraph $H$ is an ordered pair $H=(V(H),E(H))$, where $V(H)$ is a non-empty set of vertices, and $E(H)$ is a collection of non-empty subsets of $V(H)$, called hyperedges. Given an integer $k\leq|V(H)|$, a hypergraph $H$ is called $k$-uniform if every hyperedge $e\in E(H)$ has cardinality $|e|=k$. Recall that $|\cdot|$ denotes numbers of elements in the set.
@@
In random graphs and hypergraphs, the edges are random objects. For example, for each $2$-element subset of the vertex set, assign a Bernoulli random variable that equals $1$ if the subset forms an edge and 0 otherwise. Similarly, in a $k$-uniform random hypergraph, each $k$-element subset of the vertex set is assigned a Bernoulli random variable indicating whether it forms a hyperedge.
2.1.1. Erdős–Rényi Graph
Let $n$ be a natural number and let $p\in[0,1]$. The Erdős–Rényi random graph, denoted $\mathcal{G}(n,p)$, is a random graph on $n$ vertices where each edge is included independently with probability $p$ [25], [5, Chapter-10]. In other words, the Erdős–Rényi graph $\mathcal{G}(n,p)$ is a probability space $(\mathbb{G}(n),2^{\mathbb{G}(n)},{\bf P})$, where $\mathbb{G}(n)$ denotes the collection of all graphs on $n$ vertices. The probability of a particular graph $G\in\mathbb{G}(n)$ is given by:
@@ 2.1.2. $k$-uniform random hypergraphs
The $k$-uniform random hypergraph is a natural generalization of the Erdős–Rényi graphs. For any integer $k\geq 2$, the $k$-uniform random hypergraph, denoted by $H(n,k,p)$, is a random hypergraph on $n$ vertices, where each $k$-set of $n]=\{1,\ldots,n\}$ is included as a hyperedge with probability $p$ and independently. More precisely, given integers $n\geq 2$ and $1\leq k\leq n$, the $k$-uniform random hypergraph $H(n,k,p)$ is the probability space $(\mathbb{H}^{k}_{n},2^{\mathbb{H}^{k}_{n}},{\bf P})$, where $\mathbb{H}^{k}_{n}$ is the collection of all $k$-uniform hypergraphs with the vertex set $[n]$, and $2^{\mathbb{H}^{k}_{n}}$ is its power set. That is, for $H\in\mathbb{H}^{k}_{n}$ with edge set $E(H)$, the probability measure is given by
@@
2.1.3. Linial-Meshulam complex
Another well-studied generalization of the Erdős–R’enyi graph to higher dimensions is the Linial–Meshulam complex, which will appear later in our discussion. This model was introduced by Linial and Meshulam [39] for dimension $2$ and by Wallach [43] for higher dimensions.
Let $V$ be a finite set and $\mathcal{P}(V)$ denote the set of all subsets of $V$. A collection $X\subset\mathcal{P}(V)$ is said to be a simplicial complex with vertex set $V$ if $\tau\in X$ and $\sigma\subset\tau$, then $\sigma\in X$. An element of a simplicial complex is called a simplex. The dimension of a simplex $\sigma$ is defined as $|\sigma|-1$, where $|\sigma|$ denotes the number of elements in $\sigma$. An element of dimension $j$ is called a $j$-simplex. The dimension of a non-empty simplicial complex is the highest dimension among its simplices.
For a given positive integer $k\geq 2$, the Linial-Meshulam complex of dimension $k$, denoted by $Y_{k}(n,p)$, is a random simplicial complex on $n$ vertices with a complete (k1)(k-1)-dimensional skeleton and each $k$-simplices included independently with probability $p\in[0,1]$. In particular, $k=2$, the Linial Meshulam complex of dimension $1$ is the Erdős-Rényi graph.
Observe that the Linial-Meshulam complex of dimension $k$ is a random hypergraph where all simplices of dimension less than (k1)(k-1) are included as hyperedges and each $k$-dimensional simplex is included as a hyperedge with probability $p$ independently. However, it is not a $k$-uniform random hypergraph.
@@ 2.1.4. $r$-set weighted line graph
Figure 2. $H$ is a $3$-uniform hypergraph and $G_{2}(H)$ is the $2$-set line graph of $H$. Three hyperedges of $H$ correspond to three triangles, and the rest of the 22-sets are isolated vertices in this 22-set line graph.
It is worth noting that the line graph of $Y_{k}(n,p)$ is defined in [2] as the graph $(\mathcal{V}_{k},E_{n})$, where $\mathcal{V}_{k}=[\binom{n}{k}$] is the set of all (k1)(k-1)-dimensional simplices and $(\sigma,\sigma^{\prime})\in E_{n}$ if $\sigma\cup\sigma^{\prime}\in Y_{k}(n,p)$ for $\sigma,\sigma^{\prime}\in\mathcal{V}_{k}$. Observe that, the line graph of $Y_{k-1}(n,p)$ is the same as the (k1)(k-1)-set weighted line graph of $H(n,k,p)$.
@@ 2.3. Local weak convergence of hypergraphs
(4) $U_{r}(H)=U(G_{r}(H)).$
In that case, $U_{r}(H)$ is a probability measure on $(\mathbb{G}_{w}^{*},\mathcal{B}_{m})$, i.e., an element of $\mathcal{P}(\mathbb{G}_{w}^{*})$, arising from a uniformly chosen root of $H$. In that case $U_{r}(H)$ is a probability measure on $\mathcal{P}(\mathbb{G}_{w}^{*})$. Given a sequence of hypergraphs $\{H_{n}\}_{n\geq 1}$ if $\{U_{r}(H_{n})\}$ converges weakly to $\nu$, that is,
@@
then $\nu$ is called the $r$-set local weak limit of the sequence of hypergraphs $\{H_{n}\}_{n\geq 1}$.
We show that, for every $1\leq r\leq k-1$, the $r$-set local weak limit of the sequence $\{H_{n}\}$ with $H_{n}\in H(n,k,p)$ is the $d$-block Galton-Watson measure with $d=\binom{k}{r}-1$ and Poisson offspring distribution almost surely. The precise definition of the $d$-block Galton–Watson measure is given in the next subsection.
@@ 2.4.1. Galton-Watson measure
    The number of children of the root is distributed as ${\bf P}_{X}$.
@@
    Each non-root vertex is reached via an edge from its parent. If such a vertex has total degree $k$ (including the parent edge), then it has k1k-1 children, distributed according to the shifted size-biased distribution $\hat{{\bf P}}_{X}$ given by
@@
For any natural number $d$, a $d$-block Galton-Watson tree is a random graph generated by the Galton-Watson branching process in which, instead of generating a single vertex, each vertex, as offspring, produces a set of $d$ vertices called a $d$-block. Thus, if an individual in the population has a total of $n$ offspring $d$-blocks, then it has a total of $nd$ offspring vertices.
@@
    Conditioned on the number of children, the subgraphs rooted at each child (that is, each member of the child $d$-block) are independent and identically distributed copies of the $d$-block Galton–Watson graph.
@@
We now describe the rooted graphs that belong to the support of $\nu_{d,\lambda}$. The vertex set of such a rooted graph can be embedded into the set
@@
with the following ordering: $o$ is the root of the obtained graph and $o<{\bf i}$ for all ${\bf i}\neq o\in$. For $({\bf i},{\bf i}^{\prime}),({\bf j},{\bf j}^{\prime})\in\mathbb{N}^{\infty}_{d}$, suppose that $({\bf i},{\bf i}^{\prime})\in(\mathbb{N}\times[d])^{p},({\bf j},{\bf j}^{\prime})\in(\mathbb{N}\times[d])^{q}$. We set $({\bf i},{\bf i}^{\prime})<({\bf j},{\bf j}^{\prime})$ if
@@
    $p=q$, and $({\bf i},{\bf i}^{\prime})=((i_{1},i_{1}^{\prime}),\ldots,(i_{p},i_{p}^{\prime}))$, $({\bf j},{\bf j}^{\prime})=((j_{1},j_{1}^{\prime}),\ldots,(j_{p},j_{p}^{\prime}))$. Furthermore $r=\inf\{s:(i_{s},i^{\prime}_{s})\neq(j_{s},j^{\prime}_{s})\}$ with one of the followings holds:
@@
Here $Z_{n}$ represents the number of individuals in the $n$-th generation, and $X_{i,n}$ denotes the number of $d$-block offspring of the $i$-th individual in that generation. The offspring of an individual $i$ in generation $n$ are labeled by
@@
(b) The $2$-set line graph of $3$-uniform hypergraph
@@
Figure 3. A 22-set rooted hypergraph and its $2$-set line graph.
We refer the reader to [2, Section 4.2] for a detailed description of $d$-block Galton-Watson tree and the corresponding measure. Since $\nu_{d,\lambda}\in\mathcal{P}(\mathbb{G}^{*})$, that is a probability measure on the collection of all rooted graphs, $\nu_{d,\lambda}$ can be represented as a random rooted graph, we denote this random rooted graph by $(GW_{d},o)$.
For where we set $d=\binom{k}{r}-1$, the $(\binom{k}{r}-1)$-block Galton–Watson tree $(GW_{d},o)$ admits a natural representation as the $r$-set line graph of an $r$-set rooted $k$-uniform hypergraph $(T_{k},S)$, where $|S|=r$. A root $o$ corresponds to an $r$-element subset $S$ of the vertex set $V(H)$. More generally, each vertex $v$ in $GW_{d}$ corresponds to some $r$-cardinality subset $S\subset V(H)$. Every $(\binom{k}{r}-1)$-block of offspring of $v$ arises from a hyperedge $e$ of $H$ that contains $S$. Since $e$ consists of $k$ vertices, it contains exactly $\binom{k}{r}$ distinct $r$-subsets. Among these, one is $S$ itself, while the remaining $\binom{k}{r}-1$ subsets correspond to the vertices in the offspring $(\binom{k}{r}-1)$-block of $v$. Thus, $(GW_{d},o)$ is the $r$-set line graph of an $r$-set rooted $k$-uniform hypergraph $(T_{k},S)$. A $3$-uniform hypergraph $H$ and its $2$-set underlying graph is a $2$-block Galton-Watson tree illustrated in Figure 3.
2.4.3. Structure of $(T_{k},S_{r})$